A survey of 119 persons was conducted at TTC, and it was found that 89 persons carried a cell phone, 10 person carried a tablet computer, and 8 carried both cellphone and tablet.




  1. how many people carried a cellphone or a tablet?
  2. how many people carry neither a cell phone nor a tablet? 
  3. how many people carried a cell phone only?
  4. how many people carried a tablet but not a cell phone?
1
Expert's answer
2021-11-30T13:59:55-0500


1.


"N(C\cup T)=N(C)+N(T)-N(C\cap T)"

"=89+10-8=91"

2.


"N((C\cup T)')=N(U)-N(C\cup T)"

"119-91=28"

3.


"N(C-T)=N(C)-N(C\cap T)=89-8=81"

4.


"N(T-C)=N(T)-N(T\cap C)=10-8=2"